OpenStudy (anonymous):

Which of the following sets could be the sides of a right triangle? {2, 3, squareroot(13)} {5, 5, 2, squareroot(10)} {5, 12, 15}

OpenStudy (anonymous):

Top one.

OpenStudy (anonymous):

thats right

OpenStudy (sriram):

look jst by looking determine the longest side then check wheather the square root of sum of squares of other two sides= third side For eg; in the first sum longest side=13^1/2 ({2}^2 + {3}^2)^1/2= (4+9)^1/2= (13)^1/2 can do the same for others